OpenStudy (anonymous):

What is the equation of a line with a y-intercept at (0, 2) and a slope of 3?

OpenStudy (ccswims):

f(x) = mx + b

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1)

OpenStudy (midhun.madhu1987):

x1 = 0 , y1 = 2 and m = 3

OpenStudy (anonymous):

the answer is either y=2x+3, y=3x+2, y=-2x+3, or y=3x-2